OpenStudy (anonymous):

Gary teaches Native Alaskan dance. He charges a one−time processing fee of $5.00 and the cost of the classes is shown below. Let x represent the number of classes and y represent the cost of classes. Based on this information, what will it cost to take 10 classes (including processing fee)? Cost of classes (not including processing fee) Number of Classes, x 1 2 3 4 Cost of Classes, y $15.00 $27.00 $39.00 $51.00 $123.00 $125.00 $128.00 $130.00

OpenStudy (anonymous):

plot it and see what kind of function it is

OpenStudy (jdoe0001):

\(\huge a_n= a_1+(n-1)d\) d = difference between terms n= nth term, in this case 10 \(a_1\) = 1st term

OpenStudy (jdoe0001):

so, what's "d", or the difference between terms?

OpenStudy (jdoe0001):

$$ \begin{matrix} x&&y\\ \hline\\ 1&& $15\\ 2&& $27\\ 3&& $39\\ 4&& $51 \end{matrix} $$

OpenStudy (anonymous):

lol :P

OpenStudy (anonymous):

am I'm sorry but I'm really confused

OpenStudy (anonymous):

you went in reverse order it seems

OpenStudy (jdoe0001):

from the 1st class to the second, how much more do you pay?

OpenStudy (jdoe0001):

reverse? hehe, the same as typed in :)

OpenStudy (anonymous):

$13

OpenStudy (jdoe0001):

27-15 = 13?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

oh k nevermind you using a different formula. Much quicker but dont forget the formula Michele.

OpenStudy (jdoe0001):

let's check google calculator, what it says https://www.google.com/search?client=opera&rls=en&q=27+-+15&sourceid=opera&ie=utf-8&oe=utf-8&channel=suggest

OpenStudy (anonymous):

oh 12 sorry is did the math in my head it was probably off.

OpenStudy (jdoe0001):

so, how much more would you pay from the 2nd class to the 3rd class?

OpenStudy (jdoe0001):

I mean in the 3rd class you'd be able to use the hula hoop I'd think, thus is $39

OpenStudy (anonymous):

oh so i multiply 12 by 6, because 6 the 4 = 10 , then i multiply the answer 71 by $51

OpenStudy (anonymous):

but wait in the first it says there is a 5 processing fee so actully it would be15-5 27-10 = 17

OpenStudy (anonymous):

whatever, well i think I got it now so thankyou.

OpenStudy (anonymous):

it be an+5

OpenStudy (jdoe0001):

27 -15 = 12 39 - 27 = 12 51 - 39 = 12 so $$ \text{so the difference "d" between terms is} \color{red}{12}\\ \text{the term you want is the } \color{red}{10}th\\ \text{and the 1st term is }$\color{red}{15}\\ \huge a_n= a_1+(n-1)d \\ \implies \huge a_{\color{red}{10}}= \color{red}{15}+(\color{red}{10}-1)\color{red}{12} $$

OpenStudy (jdoe0001):

and as timo86m said, don't forget the "processing fee", +5 to that

OpenStudy (anonymous):

so the answer is 123